OpenStudy (anonymous):

ind the coordinates of the vertex of the parabola y = 2x2 − 4x + 9 b) One train route follows the curve y = 2x2 − 4x + 9, and the other follows y = x2. Assume that distance along both axes is measured in miles. The railroad wants to construct a north-south maintenance road PQ between the two routes. Where should the road be located so that it is as short as possible? Q(x, y) = (smaller y-value) P(x, y) = (larger y-value)

OpenStudy (anonymous):

the vertical distance between two curves is given by taking the difference between them ...\[(2x^2 - 4x + 9) - (x^2) = x^2 - 4x + 9 = (x-2)^2 + 5\]Looks like the two curves are 5 apart when x = 2 between (2,9) and (2,4)...

OpenStudy (campbell_st):

find the line of symmetry of the parabola....x = -b/2a x = -(-4)/(2x2) line of symmetry for the parabola is x = 1 substitute into th equation to find y Vertex is (1, 7)

OpenStudy (campbell_st):

the access road will be on the line x = 1/2 mid way between the two lines of symmetry. find the equation of the line between the two vertices (0,0) and (1, 7) the equation is y = 7x.... sub x = 1/2 into the equation then Q 1/2, 3 1/2)